# Job-SDF: A Multi-Granularity Dataset for Job Skill Demand Forecasting and Benchmarking

### v2

@@ Abstract
In a rapidly evolving job market, skill demand forecasting is crucial as it enables policymakers and businesses to anticipate and adapt to changes, ensuring that workforce skills align with market needs, thereby enhancing productivity and competitiveness. Additionally, by identifying emerging skill requirements, it directs individuals towards relevant training and education opportunities, promoting continuous self-learning and development. However, the absence of comprehensive datasets presents a significant challenge, impeding research and the advancement of this field. To bridge this gap, we present Job-SDF, a dataset designed to train and benchmark job-skill demand forecasting models. Based on 10.35 million public job advertisements collected from major online recruitment platforms in China between 2021 and 2023, this dataset encompasses monthly recruitment demand for 2,324 types of skills across 521 companies. Our dataset uniquely enables evaluating skill demand forecasting models at various granularities, including occupation, company, and regional levels. We benchmark a range of models on this dataset, evaluating their performance in standard scenarios, in predictions focused on lower value ranges, and in the presence of structural breaks, providing new insights for further research. Our code and dataset are publicly accessible via the https://github.com/Job-SDF/benchmark.
@@ 1 Introduction
Job skills encompass a range of abilities and competencies essential for performing tasks effectively in the workplace. These skills are broadly categorized into hard skills, such as technical and analytical abilities, and soft skills, including communication, teamwork, and adaptability [1]. Accurate forecasting of skill demand helps businesses and policymakers anticipate and address skill shortages and mismatches, and promotes skill development in high-demand areas, thereby supporting economic growth and stability [2]. By identifying emerging skill requirements, individuals are directed towards relevant training and education opportunities, fostering continuous self-learning and development to stay competitive in the labor market [3]. Moreover, forecasting informs educational and training programs, ensuring that curricula align with the labor market’s evolving needs [4].
@@
To this end, in this paper, we propose Job-SDF, a multi-granularity dataset designed for job skill demand forecasting research. Specifically, we collected 10.35 million public job advertisements from major online recruitment platforms in China, posted between January 2021 and December 2023. By extracting skill terms from job advertisement texts, we quantified the monthly skill demand at various granularities, including occupations, companies, and regions, to construct our dataset. This dataset encompasses 2,324 types of skills, 52 occupations, 521 companies, and 7 regions. We then use the Job-SDF dataset to benchmark a wide range of models for job skill demand forecasting tasks at various granularities. These models include statistical time series models (e.g., ARIMA [11]), deep learning-based methods such as RNN-based models [12, 10], Transformer-based models [13, 14, 15, 16, 17], MLP-based models [18, 19], as well as several state-of-the-art time-series forecasters [20, 21]. Performance is evaluated using regression metrics such as Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE). Additionally, we use the Symmetric Mean Absolute Percentage Error (SMAPE) [22] and Relative Root Mean Squared Error (RRMSE) [23] metrics to account for the significantly varying nature of skill demand values, which is particularly useful for evaluating model performance in predicting lower value ranges. Moreover, we further investigate the impact of structural breaks in job skill demand time series data on model performance. The Job-SDF dataset, along with data loaders, example codes for different models, and evaluation setup, are publicly available in our GitHub repository: https://github.com/Job-SDF/benchmark.
@@ 2 Related Work
Skill demand forecasting can analyze how skills evolve over time, aiding experts in evaluating technological advancements [24, 25], assessing wage inequality [26, 27], and generating employment opportunities [28]. Furthermore, the skills required in the 21st-century workplace will differ significantly from those in previous eras [29]. Predicting skill demands benefits personal career transitions and corporate management strategies.
Recently, with the rapid accumulation of data and continuous advancements in technology, skill demand forecasting has demonstrated significant vitality. Das et al. proposed a method for dynamic task allocation to investigate the evolution of job task requirements over a decade of AI innovation across different salary levels  [8]. Given the effectiveness of RNN in multi-step prediction, some researchers have integrated skill demand forecasting with RNN algorithms, achieving promising results  [9, 12]. In addition, considering the supply-demand dynamics of the labor market concurrently, CHGH designed a joint prediction model based on the encoder-decoder architecture to achieve trend prediction for both skill supply and demand sides [10]. Moreover, to capture the dynamic information of occupations, a pretraining-enhanced dynamic graph autoencoder has been developed to efficiently forecast skill demand at the occupation granularity [30].
However, the predominance of closed-source datasets has significantly elevated the barrier of researchers and constrained the pace of methodological advancements. Additionally, the current research data focuses either on macro-market skill demand predictions or on analyses at a specific granularity, neglecting multi-level labor market analysis. This limitation generally hampers the transferability of the modeling approaches.
@@ 3 Job-SDF Dataset
The Job-SDF dataset is built from job advertisements collected on online recruitment platforms, encompassing dynamic job skill demand time series data at various granularities, recorded monthly from 2021 to 2023. The dataset is CC BY-NC-SA 4.0 licensed, accessible via the URL https://github.com/Job-SDF/benchmark. We summarize the dataset construction process, task description, and dataset analysis below, with supplementary details provided in the Appendix.
@@ 3.1 Data Collection and Processing
Job Advertisement Collection. We collected public job advertisements for 52 occupations from 521 companies  [^1] Representative companies were selected based on the number of job postings they published in early 2021. on major online recruitment platforms in China, spanning from January 2021 to December 2023. After removing identical job advertisements posted simultaneously by different companies on various platforms, we obtained 10.35 million unique records. Each record contains five types of information: (1) Job Requirement, which is a text segment that outlines the specific skills required of candidates applying for the job; (2) Company, which identifies the company that posted the job advertisement; (3) Occupation, which specifies the job advertisement’s category. Our dataset encompasses 52 detailed occupations (L2-level), such as front-end development engineer and financial investment analyst. Additionally, these 52 occupations are grouped into 14 broader categories (L1-level); (4) Region, which indicates the primary geographic divisions in China where the job postings are located. These regions are classified based on their geographical orientation and include Central China, South China, North China, East China, Northeast China, Southwest China, and Northwest China; (5) Posting Time, which records the date when the job was posted, including the year, month, and day.
Job Skill Extraction. After acquiring the job advertisement data, we utilized a Named Entity Recognition (NER) model, as referenced in [31], to explicitly extract skill requirements from the Job Requirement of each advertisement.
Specifically, we first annotated a dataset for training the NER model by identifying skill terms within the job requirement texts. To achieve this, we devised a set of regular expressions tailored to the characteristics of skill descriptions and used these to match skill words in job advertisements. Subsequently, we merged all matched skill words to formulate a raw skill dictionary, including their corresponding frequencies across job advertisements. We then filtered out low-frequency words and manually annotated the raw skill dictionary to create a refined skill dictionary. Along this line, we excluded unreasonable skill words matched by the regular expressions that did not appear in the refined skill dictionary, establishing an initial correspondence between the Job Requirement and the skill requirements.
@@ 3.2 Job Skill Demand Forecasting Tasks
Our dataset includes skill demand time series data for L1-level occupations, L2-level occupations, companies, regions, and their combinations. The data spans from 2021 to 2023, with each time slice representing one month. We follow a standard protocol [32] that categorizes all time-series data into training, validation, and test sets in chronological order with a ratio of 9:1:2. In the main text, we demonstrate results with $K$ set to 6 months and consider $H$ as 3 months to evaluate the performance of different forecasting models. More settings and results can be found in the appendix and our project repository. Based on the Job-SDF dataset, other researchers can easily adjust the parameters to suit their research objectives.
@@ 3.3 Dataset Analysis
Dataset Limitation. Recent studies [35] suggest that incorporating relationships between different variants can enhance the performance of multivariate time-series forecasting. However, due to the lack of prior knowledge, our Job-SDF dataset does not yet include a graph of relationships between skills, such as predecessor-successor relationships. Instead, we constructed a skill graph based on the co-occurrence of skills in job advertisements from the training data. This graph is included in our dataset. For detailed construction methods, please refer to the Appendix.
@@ 4.1 Benchmark Models
We evaluated several SOTA time-series learning models using our proposed Job-SDF dataset. These models are categorized into six groups based on their underlying architectures: statistical time series models, RNN-based models, Transformer-based models, MLP-based models, Graph-based models, and Fourier-based models. The implementation details for each model are provided in the Appendix, and the open-source model implementations are available on our GitHub repository.
@@
RNN-based Model. RNN-based methods are effective in capturing temporal state transitions through their recurrent structures, making them widely used in various time series forecasting tasks. Notably, LSTM have demonstrated their effectiveness in predicting changes in skill shares over time [9]. However, conventional RNNs often encounter performance degradation when handling excessively long look-back windows and forecast horizons. To address this challenge, SegRNN [12] introduces segment-wise iterations, which reduce the recurrence count within RNNs, thereby significantly enhancing performance in time series forecasting tasks.
@@
Graph-based Models. Graph Neural Networks (GNNs) can learn non-Euclidean relationships, making them effective for identifying associations in structured data and generating joint representations from different perspectives. CHGH [10] uses an adaptive graph enhanced by skill co-occurrence relationships to link skill supply and demand sequences. This fusion of representations across views improves the performance of joint skill supply and demand prediction tasks. Pre-DyGAE [30] targets skill demand prediction from an occupational perspective. It builds an occupation-skill bipartite graph based on the skill demands of occupations and captures the dynamic changes in these relationships. This method allows for predicting both potential occupational skills and skill demands, leveraging a dynamic graph perspective.
@@ 5 Conclusion
In this work, we introduced Job-SDF, a dataset designed for training and benchmarking job-skill demand forecasting models. Compiled from 10.35 million public job advertisements collected from major online recruitment platforms in China between January 2021 and December 2023, this dataset includes monthly recruitment demand for 2,324 types of skills across 52 occupations, 521 companies, and 7 regions. Using this dataset, we validated a wide range of time-series forecasting approaches, including statistical models, RNN-based models, Transformer-based models, MLP-based models, Graph-based models, and Fourier-based models. Furthermore, we conducted extensive experiments to compare the performance of various methods in predicting skill demand at different granularities. We hope that Job-SDF will facilitate further research in this field.
@@ Appendix A Computational Resource
Due to inherent design and size constraints of the models combined with varying data sizes at different granularities, the deployment environments for each model are distinct. The CHGH model, which requires over 80GB of memory, is exclusively deployed on CPU platforms to accommodate its substantial resource demands. In contrast, the PreDyGAE model operates solely on GPU infrastructure, leveraging the computational efficiencies of the NVIDIA A800 GPUs. For other models, deployment strategies are tailored according to the granularity of the data. Experiments at the labor market, regions, L1 occupations, L2 occupations, and Region & L1 occupations granularities are conducted on GPUs, capitalizing on the enhanced processing capabilities of these units for handling moderate data volumes. However, at the granularities of Region & L2 and company, where data volumes are significantly larger, deployment shifts to CPUs. Overall, the training time of different models are shown in Table S5.
Table S5: Training time (minute) of different models for job skill demand forecsting.
@@ B.1 Repeated Experiments on Job Skill Demand Forecasting
Table S6 displays the mean and standard deviation results of repeated experiments on the benchmark models for the job skill demand forecasting task as presented in the main text. Initially, we supplement the experimental results at the overall labor market level and regional granularity, where the RMSE averages over 1000. In cases of coarser granularity, due to the larger base of demand values, the prediction deviations are significant.
Table S6: Overall performance comparisons on repeated experiments.
@@
Table S7: Performance comparisons on skill demand series with structural breaks.
@@
Table S7 presents the results of repeated experiments on forecasting skill demand sequences that have undergone structural breaks. Initially, the overall errors are quite pronounced, underscoring the challenge of accurately predicting these skills. Moreover, FiLM performs well on most metrics, which further verifies the robustness of FiLM.
@@ B.2 Job Skill Demand Forecasting for Low-Frequency Skills
In multigranular skill demand sequences, a significant number of skills remain inactive or in low frequency over extended periods. These skills might continue to have low demand in the future (indicating low importance), or they might suddenly gain interest from certain professions or companies, leading to rapid growth. In this study, we define low-frequency skills as those that appear fewer than twice in the time slices of the training set. Predicting the demand for these skills is challenging because their data points are predominantly zero during training, resulting in a lack of effective observational data. Therefore, we specifically present the demand prediction results of the existing benchmark models for these low-frequency skills, which could serve as a potential direction for further research.
@@ Results.
We continued to test the demand prediction effect on low-frequency skills using the benchmark models described in the main text, and the results are shown in Table S8. From this, we can draw the following conclusions: Firstly, there is a significant increase in the error on the RRMSE metric, indicating that low-demand skills are difficult to predict accurately. Secondly, Koopa has the best predictive performance in this scenario. We also found that the performance of SegRNN significantly decreases, suggesting that SegRNN’s segment learning approach is not suitable for predicting low-frequency skill demands due to a lack of effective observational data, rendering the learning segments meaningless.
Table S8: Performance comparisons on skill demand series with low-frequency.
@@
B.3 Skill Co-occurrence Graph Enhanced Job Skill Demand Forecasting
Benchmark Models
Results
We have implemented a series of graph-based multivariate time series forecasting methods based on the co-occurrence graph and verified their experimental effects under the three scenarios discussed above. Firstly, Table S9 presents the overall performance of the methods based on the co-occurrence graph for skill demand forecasting. It is observed that the prediction accuracy significantly declines across the overall labor market. However, as the granularity of the forecast becomes finer, the model performance improves, and at a finer granularity, the EvolveGCN method outperforms the state-of-the-art (SOTA) methods mentioned in the main text considerably. We analyze that the finer the granularity, the more accurately the co-occurrence graph reflects the associations between skills, while coarser granularity might introduce excessive noise leading to decreased model performance. The fine-grained co-occurrence graph accurately reflects the interrelationships between skills at different granularities, which aids in enhancing the model’s prediction accuracy. Secondly, we find significant differences in the RRMSE metric among these methods, with EvolveGCN showing superior performance because it can learn the evolution of GCN parameter weights over time, thus capturing the evolving dependencies among edges. Therefore, based on the provided co-occurrence graph, it can effectively learn the evolution of skill relationships, which is beneficial for dynamic prediction of skill demand.
For skill demand forecasting in scenarios involving structural breaks, as shown in Table S10, the improvements in the methods based on the co-occurrence graph are greater than those in the overall skill demand forecasting task. This suggests that skills experiencing structural breaks have strong interconnections, and the co-occurrence graph helps the model to identify the patterns of skill demand sequences that are likely to undergo structural breaks, thus further enhancing the prediction effectiveness for this category of skills.
In the task of predicting low-frequency skills, as shown in Table S11, methods like GConvLSTM significantly outperform EvolveGCN. This is due to the sparse observable data for these skills, which leads to sparse connectivity edges on the co-occurrence graph.
Table S9: Performance comparisons on skill demand series with GNN-based methods.
@@ Results
Table S10: Performance comparisons on skill demand series with structural breaks with GNN-based methods.
@@
Table S11: Performance comparisons on low-frequency skill demand series with GNN-based methods.
@@
B.4 Job Skill Demand Proportion Forecasting
Results
We continue to utilize the benchmark models described in the main text for this task, and the results, as shown in Table S12, lead to the following conclusions: Firstly, the best-performing model on the task of forecasting the proportion of skill demand is Koopa. This model, integrating time series decomposition and Fourier transformations, effectively captures the distribution changes in demand proportions. Secondly, there is a significant variation in performance across models in this task. For example, models like DLinear perform poorly on this task, though they are reasonably effective in skill demand forecasting. We analyze that predicting percentages is distinct from forecasting skill demand, as percentage predictions are also influenced by the demand for other skills at the same granularity. Therefore, simple linear models are not advantageous for capturing the complex interrelations and influences among multiple pieces of information.
Table S12: Performance comparisons on skill demand proportion forecasting.

### v3

@@ Abstract
In a rapidly evolving job market, skill demand forecasting is crucial as it enables policymakers and businesses to anticipate and adapt to changes, ensuring that workforce skills align with market needs, thereby enhancing productivity and competitiveness. Additionally, by identifying emerging skill requirements, it directs individuals towards relevant training and education opportunities, promoting continuous self-learning and development. However, the absence of comprehensive datasets presents a significant challenge, impeding research and the advancement of this field. To bridge this gap, we present Job-SDF, a dataset designed to train and benchmark job-skill demand forecasting models. Based on millions of public job advertisements collected from online recruitment platforms, this dataset encompasses monthly recruitment demand. Our dataset uniquely enables evaluating skill demand forecasting models at various granularities, including occupation, company, and regional levels. We benchmark a range of models on this dataset, evaluating their performance in standard scenarios, in predictions focused on lower value ranges, and in the presence of structural breaks, providing new insights for further research. Our code and dataset are publicly accessible via the https://github.com/Job-SDF/benchmark.
@@ 1 Introduction
Job skills encompass a range of abilities and competencies essential for performing tasks effectively in the workplace. These skills are broadly categorized into hard skills, such as technical and analytical abilities, and soft skills, including communication, teamwork, and adaptability [1]. Accurate forecasting of skill demand helps businesses and policymakers anticipate and address skill shortages and mismatches, and promotes skill development in high-demand areas, thereby supporting economic growth and stability [2, 3]. By identifying emerging skill requirements, individuals are directed towards relevant training and education opportunities, fostering continuous self-learning and development to stay competitive in the labor market [4, 5, 6, 7, 8, 9, 10]. By tracking skill demand trends, employers gain deeper insight into recruits’ priorities, enhancing person-job fit. [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. Moreover, forecasting informs educational and training programs, ensuring that curricula align with the labor market’s evolving needs [22, 23, 24].
@@
To this end, in this paper, we propose Job-SDF, a multi-granularity dataset designed for job skill demand forecasting research. Specifically, we collected millions of public job advertisements from online recruitment platforms. By extracting skill terms from job advertisement texts, we quantified the monthly skill demand at various granularities, including occupations, companies, and regions, to construct our dataset. This dataset encompasses 2,324 types of skills, 52 occupations, 521 companies, and 7 regions. We then use the Job-SDF dataset to benchmark a wide range of models for job skill demand forecasting tasks at various granularities. These models include statistical time series models (e.g., ARIMA [31]), deep learning-based methods such as RNN-based models [32, 30], Transformer-based models [33, 34, 35, 36, 37], MLP-based models [38, 39], as well as several state-of-the-art time-series forecasters [40, 41]. Performance is evaluated using regression metrics such as Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE). Additionally, we use the Symmetric Mean Absolute Percentage Error (SMAPE) [42] and Relative Root Mean Squared Error (RRMSE) [43] metrics to account for the significantly varying nature of skill demand values, which is particularly useful for evaluating model performance in predicting lower value ranges. Moreover, we further investigate the impact of structural breaks in job skill demand time series data on model performance. The Job-SDF dataset, along with data loaders, example codes for different models, and evaluation setup, are publicly available in our GitHub repository: https://github.com/Job-SDF/benchmark.
@@ 2 Related Work
Skill demand forecasting can analyze how skills evolve over time, aiding experts in evaluating technological advancements [44, 45, 46], assessing wage inequality [47, 48, 49], and generating employment opportunities [50]. Furthermore, the skills required in the 21st-century workplace will differ significantly from those in previous eras [51]. Predicting skill demands benefits personal career transitions and corporate management strategies.
Recently, with the rapid accumulation of data and continuous advancements in technology, skill demand forecasting has demonstrated significant vitality. Das et al. proposed a method for dynamic task allocation to investigate the evolution of job task requirements over a decade of AI innovation across different salary levels  [28]. Given the effectiveness of RNN in multi-step prediction, some researchers have integrated skill demand forecasting with RNN algorithms, achieving promising results  [29, 32]. In addition, considering the supply-demand dynamics of the labor market concurrently, CHGH designed a joint prediction model based on the encoder-decoder architecture to achieve trend prediction for both skill supply and demand sides [30]. Moreover, to capture the dynamic information of occupations, a pretraining-enhanced dynamic graph autoencoder has been developed to efficiently forecast skill demand at the occupational granularity [52].
However, the predominance of closed-source datasets has significantly elevated the barrier of researchers and constrained the pace of methodological advancements. While open-source skill-related datasets such as O*NET [53] and ESCO [54] provide skill taxonomies, they do not quantify skill demand. Furthermore, the current research data focuses either on macro-market skill demand predictions or analyses at a specific granularity, neglecting multi-level labor market analysis. This limitation generally hampers the transferability of the modeling approaches.
@@ 3 Job-SDF Dataset
The Job-SDF dataset is built from job advertisements collected on online recruitment platforms, encompassing dynamic job skill demand time series data at various granularities, recorded monthly. The dataset is CC BY-NC-SA 4.0 licensed, accessible via the URL https://github.com/Job-SDF/benchmark. We summarize the dataset construction process, task description, and dataset analysis below.
@@ 3.1 Data Collection and Processing
Job Advertisement Collection. We collected public job advertisements for 52 occupations from 521 companies on online recruitment platforms. We obtained unique records after removing identical job advertisements posted simultaneously by different companies on various platforms. Each record contains five types of information: (1) Job Requirement, which is a text segment that outlines the specific skills required of candidates applying for the job; (2) Company, which identifies the company that posted the job advertisement; (3) Occupation, which specifies the job advertisement’s category. Our dataset encompasses 52 detailed occupations (L2-level), such as front-end development engineer and financial investment analyst. Additionally, these 52 occupations are grouped into 14 broader categories (L1-level); (4) Region, which indicates the primary geographic divisions in China where the job postings are located. These regions are classified based on their geographical orientation; (5) Posting Time, which records the date when the job was posted, including the year, month, and day.
Job Skill Extraction. After acquiring the job advertisement data, we utilized a Named Entity Recognition (NER) model, as referenced in [55, 56, 57, 58], to explicitly extract skill requirements from the Job Requirement of each advertisement. Specifically, we first annotated a dataset for training the NER model by identifying skill terms within the job requirement texts. To achieve this, we devised a set of regular expressions tailored to the characteristics of skill descriptions and used these to match skill words in job advertisements. Subsequently, we merged all matched skill words to formulate a raw skill dictionary, including their corresponding frequencies across job advertisements. We then filtered out low-frequency words and manually annotated the raw skill dictionary to create a refined skill dictionary. Along this line, we excluded unreasonable skill words matched by the regular expressions that did not appear in the refined skill dictionary, establishing an initial correspondence between the Job Requirement and the skill requirements.
@@ 3.2 Job Skill Demand Forecasting Tasks
Our dataset includes skill demand time series data for L1-level occupations, L2-level occupations, companies, regions, and their combinations. We follow a standard protocol [59] that categorizes all time-series data into training, validation, and test sets in chronological order with a ratio of 9:1:2. In the main text, we demonstrate results with $K$ set to 6 months and consider $H$ as 3 months to evaluate the performance of different forecasting models. More settings and results can be found in the Appendix B and our project repository. Based on the Job-SDF dataset, other researchers can easily adjust the parameters to suit their research objectives.
@@ 3.3 Dataset Analysis
Dataset Limitation. Recent studies [62] suggest that incorporating relationships between different variants can enhance the performance of multivariate time-series forecasting. However, due to the lack of prior knowledge, our Job-SDF dataset does not yet include a graph of relationships between skills, such as predecessor-successor relationships. Instead, we constructed a skill graph based on the co-occurrence of skills in job advertisements from the training data. This graph is included in our dataset.
@@ 4.1 Benchmark Models
We evaluated several SOTA time-series learning models using our proposed Job-SDF dataset. These models are categorized into six groups based on their underlying architectures: statistical time series models, RNN-based models, Transformer-based models, MLP-based models, Graph-based models, and Fourier-based models. The implementation details for each model are provided in the Appendix B, and the open-source model implementations are available on our GitHub repository.
@@
RNN-based Model. RNN-based methods are effective in capturing temporal state transitions through their recurrent structures, making them widely used in various time series forecasting tasks [65, 66, 67, 68, 69]. Notably, LSTM have demonstrated their effectiveness in predicting changes in skill shares over time [29]. However, conventional RNNs often encounter performance degradation when handling excessively long look-back windows and forecast horizons. To address this challenge, SegRNN [32] introduces segment-wise iterations, which reduce the recurrence count within RNNs, thereby significantly enhancing performance in time series forecasting tasks.
@@
Graph-based Models. Graph Neural Networks (GNNs) can learn non-Euclidean relationships, making them effective for identifying associations in structured data and generating joint representations from different perspectives [73, 74, 75, 76]. CHGH [30] uses an adaptive graph enhanced by skill co-occurrence relationships to link skill supply and demand sequences. This fusion of representations across views improves the performance of joint skill supply and demand prediction tasks. Pre-DyGAE [52] targets skill demand prediction from an occupational perspective. It builds an occupation-skill bipartite graph based on the skill demands of occupations and captures the dynamic changes in these relationships. This method allows for predicting both potential occupational skills and skill demands, leveraging a dynamic graph perspective.
@@ 5 Conclusion
In this work, we introduced Job-SDF, a dataset designed for training and benchmarking job-skill demand forecasting models. Compiled from millions of public job advertisements collected from online recruitment platforms, this dataset includes monthly recruitment demand for 2,324 types of skills across 52 occupations, 521 companies, and 7 regions. Using this dataset, we validated a wide range of time-series forecasting approaches, including statistical models, RNN-based models, Transformer-based models, MLP-based models, Graph-based models, and Fourier-based models. Furthermore, we conducted extensive experiments to compare the performance of various methods in predicting skill demand at different granularities. We hope that Job-SDF will facilitate further research in this field.
Acknowledgements
This work was supported in part by the National Key R&D Program of China (Grant No.2023YFF0725001), in part by the National Natural Science Foundation of China (Grant No.92370204), in part by the guangdong Basic and Applied Basic Research Foundation (Grant No.2023B1515120057), in part by Guangzhou-HKUST (GZ) Joint Funding Program (Grant No.2023A03J0008), Education Bureau of Guangzhou Municipality, in part by Nansha Postdoctoral Research Project, and in part by the National Natural Science Foundation of China (Grant No.62176014), the Fundamental Research Funds for the Central Universities.
@@ Appendix A Computational Resource
Due to inherent design and size constraints of the models combined with varying data sizes at different granularities, the deployment environments for each model are distinct. The CHGH model, which requires over 80GB of memory, is exclusively deployed on CPU platforms to accommodate its substantial resource demands. In contrast, the PreDyGAE model operates solely on GPU infrastructure, leveraging the computational efficiencies of the NVIDIA A800 GPUs. For other models, deployment strategies are tailored according to the granularity of the data. Experiments at the labor market, regions, L1 occupations, L2 occupations, and Region & L1 occupations granularities are conducted on GPUs, capitalizing on the enhanced processing capabilities of these units for handling moderate data volumes. However, at the granularities of Region & L2 and company, where data volumes are significantly larger, deployment shifts to CPUs. Overall, the training time of different models are shown in Table 5.
Table 5: Training time (minute) of different models for job skill demand forecsting.
@@ B.1 Repeated Experiments on Job Skill Demand Forecasting
Table 6 displays the mean and standard deviation results of repeated experiments on the benchmark models for the job skill demand forecasting task as presented in the main text. Initially, we supplement the experimental results at the overall labor market level and regional granularity, where the RMSE averages over 1000. In cases of coarser granularity, due to the larger base of demand values, the prediction deviations are significant.
Table 6: Overall performance comparisons on repeated experiments.
@@
Table 7: Performance comparisons on skill demand series with structural breaks.
@@
Table 7 presents the results of repeated experiments on forecasting skill demand sequences that have undergone structural breaks. Initially, the overall errors are quite pronounced, underscoring the challenge of accurately predicting these skills. Moreover, FiLM performs well on most metrics, which further verifies its robustness.
@@ B.2 Job Skill Demand Forecasting for Low-Frequency Skills
In multigranular skill demand sequences, a significant number of skills remain inactive or in low frequency over extended periods. These skills might continue to have low demand in the future (indicating low importance), or they might suddenly gain interest from certain professions or companies, leading to rapid growth. In this study, we define low-frequency skills as those that appear fewer than twice in the time slices of the training set. Predicting the demand for these skills is challenging because their data points are predominantly zero during training, resulting in a lack of effective observational data. Therefore, we specifically present the demand prediction results of the existing benchmark models for these low-frequency skills.
@@ Results.
We continued to test the demand prediction effect on low-frequency skills using the benchmark models described in the main text, and the results are shown in Table 8. From this, we can draw the following conclusions: Firstly, there is a significant increase in the error on the RRMSE metric, indicating that low-demand skills are difficult to predict accurately. Secondly, Koopa has the best predictive performance in this scenario. We also found that the performance of SegRNN significantly decreases, suggesting that SegRNN’s segment learning approach is not suitable for predicting low-frequency skill demands due to a lack of effective observational data, rendering the learning segments meaningless.
Table 8: Performance comparisons on skill demand series with low-frequency.
@@
Table 9: Performance comparisons on skill demand series with GNN-based methods.
@@
Table 10: Performance comparisons on skill demand series with structural breaks with GNN-based methods.
@@
Table 11: Performance comparisons on low-frequency skill demand series with GNN-based methods.
@@
Table 12: Performance comparisons on skill demand proportion forecasting.
@@
B.3 Skill Co-occurrence Graph Enhanced Job Skill Demand Forecasting
Benchmark Models
Results
We have implemented a series of graph-based multivariate time series forecasting methods based on the co-occurrence graph and verified their experimental effects under the three scenarios discussed above. Firstly, Table 9 presents the overall performance of the methods based on the co-occurrence graph for skill demand forecasting. It is observed that the prediction accuracy significantly declines across the overall labor market. However, as the granularity of the forecast becomes finer, the model performance improves, and at a finer granularity, the EvolveGCN method outperforms the state-of-the-art (SOTA) methods mentioned in the main text considerably. We analyze that the finer the granularity, the more accurately the co-occurrence graph reflects the associations between skills, while coarser granularity might introduce excessive noise leading to decreased model performance. The fine-grained co-occurrence graph accurately reflects the interrelationships between skills at different granularities, which aids in enhancing the model’s prediction accuracy. Secondly, we find significant differences in the RRMSE metric among these methods, with EvolveGCN showing superior performance because it can learn the evolution of GCN parameter weights over time, thus capturing the evolving dependencies among edges. Therefore, based on the provided co-occurrence graph, it can effectively learn the evolution of skill relationships, which is beneficial for dynamic prediction of skill demand.
For skill demand forecasting in scenarios involving structural breaks, as shown in Table 10, the improvements in the methods based on the co-occurrence graph are greater than those in the overall skill demand forecasting task. This suggests that skills experiencing structural breaks have strong interconnections, and the co-occurrence graph helps the model to identify the patterns of skill demand sequences that are likely to undergo structural breaks, thus further enhancing the prediction effectiveness for this category of skills.
In the task of predicting low-frequency skills, as shown in Table 11, methods like GConvLSTM significantly outperform EvolveGCN. This is due to the sparse observable data for these skills, which leads to sparse connectivity edges on the co-occurrence graph.
B.4 Job Skill Demand Proportion Forecasting
Results
We continue to utilize the benchmark models described in the main text for this task, and the results, as shown in Table 12, lead to the following conclusions: Firstly, the best-performing model on the task of forecasting the proportion of skill demand is Koopa. This model, integrating time series decomposition and Fourier transformations, effectively captures the distribution changes in demand proportions. Secondly, there is a significant variation in performance across models in this task. For example, models like DLinear perform poorly on this task, though they are reasonably effective in skill demand forecasting. We analyze that predicting percentages is distinct from forecasting skill demand, as percentage predictions are also influenced by the demand for other skills at the same granularity. Therefore, simple linear models are not advantageous for capturing the complex interrelations and influences among multiple pieces of information.